OpenStudy (anonymous):

I Need Help Subtracting Fractions With No Common Denominatoer????(:

OpenStudy (anonymous):

Do you have an example or would you like one to practice?

OpenStudy (anonymous):

Yes 4/5 subtract 1/3

myininaya (myininaya):

\[\frac{a}{c}-\frac{b}{d}=\frac{ad-bc}{cd}\]

OpenStudy (anonymous):

I Dont Understand That?

myininaya (myininaya):

replace all the a's you see with 4 replace all the c's you see with 5 replace all the b's you see with 1 replace all the d's you see with 3

OpenStudy (anonymous):

Okay \[\frac{4}{5}-\frac{1}{3}=\frac{4*3}{5*3}-\frac{1*5}{3*5}\] \[\frac{12}{15}-\frac{5}{15}=\frac{7}{15}\]

myininaya (myininaya):

that works too

OpenStudy (anonymous):

Thank You So Much!(:

OpenStudy (anonymous):

Can Yall Hep Me With Another Problem? 5/8 subtract 1/6?